OpenStudy (anonymous):

Find the exact circumference, in terms of , of a circle with the given radius. x/2 miles C = miles

OpenStudy (anonymous):

in terms of pie

OpenStudy (anonymous):

c = 2 Pi r if r=x/2, then what do you thinnk c=?

OpenStudy (anonymous):

x/4

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2*pi*x/2=pi*x

OpenStudy (anonymous):

2 times 3.14 times /2 = 3.14 times x

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

soyour answer is 3.14x

OpenStudy (anonymous):

yes or more precisely pi*x

OpenStudy (anonymous):

Find the approximate value of the circumference of a circle with the given radius. Use = 3.14. Round your answer to the nearest tenth. 4 inches C = inches

OpenStudy (anonymous):

so if i use the formula you told me 2*pi*x/2=pi*x where would i substitute the 4 at

OpenStudy (anonymous):

c= pi* 2r

OpenStudy (anonymous):

c = 3.14x 2r would be 16 3.14x16

OpenStudy (anonymous):

correct